OpenStudy (kkutie7):

water leaks out of a barrel at a rate proportional to the square root of the depth, D, of the water at that time. Depth is measure in inches and time i measured in hours. (don't worry about units my professor doesn't care) There are three parts I'll post below.

OpenStudy (kkutie7):

(a) Write a differential equation for the depth of water at time t. Assume the proportionality constant is k>0. \[\frac{dD}{dt}=-k \sqrt{D} \]

OpenStudy (kkutie7):

(b) If the depth of the water is initially 36 inches, find D(t). Your answer will have k in it. \[\int D^{-\frac{1}{2}}dD=\int-kdt\] \[2D^{\frac{1}{2}}=-kt+C\] \[D^{\frac{1}{2}}=C-\frac{kt}{2}\] \[D=C-(\frac{kt}{2})^{2}\] \[36=C \rightarrow D=36-(\frac{kt}{2})^{2}\] got it wrong =(

OpenStudy (kkutie7):

I'm struggling to see my error

zepdrix (zepdrix):

Oh, you silly billy. \[\large\rm \left[C-\frac{kt}{2}\right]^2\quad\ne\quad C^2-\left(\frac{kt}{2}\right)^2\]

zepdrix (zepdrix):

Remember that old FOIL business? ;)

OpenStudy (kkutie7):

darn I forget to write the power of two

zepdrix (zepdrix):

I hope you understand what I'm saying. The fact that you absorbed the square into the C is not the problem.

zepdrix (zepdrix):

The problem is: You should end up with `three terms` when you expand out a 2nd degree binomial. Not just two terms.

OpenStudy (kkutie7):

no I get it i separated them when I shouldn't have

zepdrix (zepdrix):

oh ok c:

zepdrix (zepdrix):

Yah I would input the initial data before squaring. The algebra works out a little easier that way. Either is fine though. You're inputting this to a website or something?

OpenStudy (kkutie7):

yup

OpenStudy (kkutie7):

so this way you get \[36-6kt+(\frac{kt}{2})^{2}\]

zepdrix (zepdrix):

If you like to expand out the square :) yes! I think it looks a tad nicer in this form though,\[\large\rm D(t)=\frac{1}{4}\left(12-kt\right)^2\]But yay I think you got it \c:/

OpenStudy (kkutie7):

Yes! thanks. I'm gonna continue to work out the rest of the problem on here just in case I goof again =)

OpenStudy (kkutie7):

(c) If the depth of the water drops to 35 inches in 1 hour, how long will it take for the water to leak out of the bucket. \[35=\frac{1}{4}(12-k(1))^{2}\] \[\sqrt{140}=12-k\] \[12-\sqrt{140}=k =0.16784043\] \[D=\frac{1}{4}(12-t(12-\sqrt{140}))^{2}\] \[0=\frac{1}{4}(12-12t+t\sqrt{140}))^{2}\] \[0=(12-12t+t\sqrt{140})^{2}\] \[0=(284-24\sqrt{140})t^{2}+(24\sqrt{140}-288)t+144\]

OpenStudy (kkutie7):

\[t=6(6+\sqrt{35})hours\]

OpenStudy (kkutie7):

71.5 hours

zepdrix (zepdrix):

Hmm I would probably square root each side from this step:\[0=(12-12t+t\sqrt{140})^{2}\]Instead of squaring it out >.<

OpenStudy (kkutie7):

oh I know but i had to show set by step with squaring as I entered it... what a drain

zepdrix (zepdrix):

oh haha